# ScarFinder: a detector of optimal scar trajectories in quantum many-body dynamics

### v1

@@ V Identifying stable islands in Poincaré sections
In few-body dynamical systems, periodic dynamics and atypical eigenstates can emerge from the coexistence of regular and chaotic regions, resulting from weakly-broken integrability described by the Kolmogorov-Arnold-Moser (KAM) theorem [86, 87]. This phenomenon is known as mixed phase space and it has many physical realizations, including the Chirikov map, the Fermi-Pasta-Ulam system of coupled oscillators, the Lorentz attractor, and many other [88]. For a general dynamical system, mixed phase space is diagnosed by recording successive intersections of trajectories with a chosen lower-dimensional transverse surface in phase space – a map known as the Poincaré section [89]. Periodic trajectories correspond to stationary points of the Poincaré map, while during the chaotic dynamics, the system returns to the same surface at a location that is generally far away from a previous encounter.
Ref. [71] demonstrated that analogues of Poincaré sections can be identified in many-body quantum systems, once their dynamics is projected to a low-dimensional MPS manifold. The resulting TDVP equations of motion within the manifold then serve as an effective mapping of the many-body quantum system onto a few-body classical (non-linear) dynamical system. For the latter, standard tools [89] can be employed to obtain Poincaré sections. However, such analyses become increasingly cumbersome in higher-dimensional manifolds. A pertinent practical question is: can one directly identify the centers of KAM tori associated with scar dynamics?
From the above description of the Poincaré section, it should be clear that it has many things in common with the ScarFinder algorithm. As discussed in Sec. II, the ScarFinder algorithm can be viewed as generating discrete-time dynamics on the variational manifold. Although the theoretical justification of ScarFinder naturally favors larger projection time steps $\Delta t$, the method remains well-defined in the continuous-time limit $\Delta t\rightarrow 0$. In this limit, ScarFinder dynamics become equivalent to the TDVP evolution equations [71]. This connection implies that, for a given MPS ansatz describing the manifold $\mathcal{M}$, the continuous-time limit $\Delta t\rightarrow 0$ of the ScarFinder algorithm should recover precisely the TDVP dynamics within $\mathcal{M}$. Since any numerical implementation of ScarFinder necessarily uses finite time steps $\Delta t$, such discretizations will inevitably introduce small, systematic deviations from the ideal TDVP trajectory. Surprisingly, rather than causing the dynamics to become chaotic, these deviations accumulate constructively, driving the trajectory towards stable fixed points—precisely the centers of stable islands in the mixed phase space portrait. This motivates the following generalization of the ScarFinder algorithm for exploring Poincaŕe sections.
@@ V.1 ScarFinder algorithm for stable islands in Poincaré sections
Input: An appropriate manifold $\mathcal{M}$, Hamiltonian $\hat{H}$, ScarFinder projection time step $\Delta t$, ScarFinder iteration steps $N_{\text{step}}$, Poincaré sampling number $N_{\text{sample}}$.
@@
4:         Evolve $|\psi\rangle$ to $|\psi^{\prime}\rangle=\exp(-i\hat{H}\Delta t)|\psi\rangle$;
5:          Project $|\psi^{\prime}\rangle$ back onto $\mathcal{M}$ by maximizing the logarithmic overlap, $\max_{A({\bm{\theta}})}\mathcal{N}[\psi^{\prime},\psi(A(\bm{\theta}))]$, $|\psi^{\prime}\rangle\to\hat{P}_{\mathcal{M}}|\psi^{\prime}\rangle$;
@@
9:         Update $|\psi\rangle\to|\psi^{\prime}\rangle$.
@@
We note that, in general, the projection step (line 5) should include an energy correction to ensure that the target energy density is preserved. This can be achieved by slightly modifying the projection procedure to incorporate an energy constraint during the maximization. However, for certain models such as the PXP model studied in Sec. V.2 below, this correction turns out to be unnecessary, as the chosen variational manifold inherently preserves the energy.
@@ V.2 Application to PXP mixed phase space
Figure 6: (a)-(b) Comparison of Poincaré sections obtained via TDVP and ScarFinder. Trajectories obtained from TDVP (blue points) and ScarFinder (orange points with numbers indicating iteration steps), starting from the same initial condition. The trajectory under ScarFinder spirals rapidly towards the center of the stable island, corresponding exactly to the robust scar trajectory identified as the “*-state” in Ref. [71]. (c) Our reproduction of the TDVP Poincaré section from Ref. [71], displaying a mixture of stable islands and chaotic regions. (d) ScarFinder directly identifies the centers of stable islands in (c).
Let us illustrate the Poincaré-ScarFinder algorithm on the example considered in Ref. [71]. We assume the PXP model with a simple two-dimensional manifold defined by the following MPS ansatz, first introduced in Ref. [53]:
@@
Figures 6(a)-(b) illustrate the contrast between TDVP and ScarFinder dynamics. While TDVP trajectories trace invariant tori, the finite-step ScarFinder trajectory spirals to a unique attractor, located precisely at the center of a stable island. In Fig. 6(c)-(d), we present additional examples, confirming that ScarFinder systematically identifies the main stable island centers, provided that a sufficient number of initial conditions are sampled.
@@
For smaller islands, there are small, athough visible, deviations between TDVP island centers and fixed points identified by ScarFinder. Here we analyze these differences in detail. By extensively sampling 640 random initial states within the three-site unit-cell MPS manifold for the PXP model, Eq. (19), we identify a total of 16 distinct fixed-point (within $\theta_{1}\in[0.5\pi,1.5\pi],\ \theta_{3}\in[0,\pi]$) solutions using the ScarFinder algorithm, as shown in Fig. 7(a). We find that the fixed points labeled “1” and “2” correspond exactly (within numerical precision) to the “*-state” previously discovered in Ref. [71]. These points coincide precisely with the centers of the largest stable islands identified from the TDVP-generated Poincaré sections.
@@
Figure 7: (a) Distribution of 16 distinct fixed-point solutions identified by the ScarFinder algorithm from a total of 640 random initial state samplings within the MPS manifold given by Eq. (19). We label the fixed-point sets according to the frequency of getting that solution. That red star in the figure marks the position of the optimized state from TDVP-Poincaré section method. The red circle highlights two points from different methods, illustrating the dynamics in panels (b) and (c). (b) Comparison of bipartite entanglement entropy dynamics starting from optimized initial states obtained via TDVP Poincaré section (blue) and ScarFinder (orange line). (c) Fidelity revival dynamics for the two optimized scar initial states. The state found by ScarFinder exhibits slightly enhanced revivals.
[…]
@@
We thus conclude that the ScarFinder algorithm not only efficiently locates optimal scar states, but in certain cases, even provides better initial conditions compared to states obtained from direct exploration of TDVP-generated mixed-phase spaces. This can be understood by noting that TDVP restricts quantum dynamics to a particular variational ansatz, unlike ScarFinder which utilizes information from exact quantum dynamics during each iteration. The deviation between TDVP and the exact quantum dynamics is characterized using “quantum leakage” [54]. Since scarred trajectories are typically found in regions of small leakage, there is a strong correspondence between their TDVP trajectory and exact quantum dynamics. However, even in these cases the leakage is non-zero and some relevant information has been lost, therefore it is unsurprising that ScarFinder, can improve upon these results as it does not suffer from any leakage effects.
@@ VI Conclusions
Beyond reproducing the known QMBS phenomenology, our application of ScarFinder to the PXP model gave several intriguing new results. In particular, our observations in Fig. 4 and Fig. 5 hint at an equal energy spacing between (some) scar eigenstates in the thermodynamic limit. Previously, the scar dynamics observed from the $|\mathbb{Z}_{2}\rangle$ state were attributed to an approximately equal energy spacing among these eigenstates [15, 54]. However, since the $|\mathbb{Z}_{2}\rangle$ state does not exhibit perfect revivals and exact diagonalization is limited by finite-size effects, it has remained unclear whether the equal energy spacing persists in the thermodynamic limit or if it is merely an approximation. Our improved superposition state in the form of an iMPS and the trend in Figs. 4-5 suggest that equal energy spacing is an exact property of some eigenstates in the middle of the spectrum of the PXP model in the thermodynamic limit. Indeed, this is also in line with recent work [60] which found several exact MPS eigenstates in the spectrum of the PXP model.
@@
We emphasize that the utility of ScarFinder is not limited to QMBS models and it can also aid the identification and understanding of other types of atypical dynamics in chaotic models. For example, in the paradigmatic quantum Ising model in the presence of both transverse and longitudinal fields, Ref. [74] found certain product states that undergo “weak” thermalization in contrast to the majority of other initial states that strongly thermalize. While these states were subsequently related to low-lying quasiparticle excitations [75], their existence is nonetheless surprising given that the model is believed to be “fully” chaotic and obey the ETH. In Appendix D, we show that ScarFinder successfully captures these states and allows to “refine” them by including entanglement. More importantly, ScarFinder allows to search for weakly-thermalizing states, like the ones in Ref. [74], systematically and in general models, without prior knowledge about their underlying physics.
We conclude by mentioning several noteworthy advantages of ScarFinder. In 1D systems, the algorithm can be straightforwardly implemented using MPS states, making it numerically efficient. Moreover, the algorithm can directly work in the thermodynamic limit, minimizing finite-size effects. Remarkably, finite discretization errors, rather than compromising stability of the algorithm, conspire to guide trajectories toward optimal scar states, providing a method for discovering and optimizing QMBSs. On the other hand, we showed that ScarFinder allows to directly identify the centers of stable islands in Poincaré sections, removing the need for complicated visualizations of high-dimensional phase spaces. These technical advantages make ScarFinder a versatile tool for identifying non-thermal behaviors in many-body systems, in particular in the field of quantum simulation and designing long-lived quantum states.
@@
Note Added: During the completion of this work, we became aware of a related study by Petrova et al. [90], where a different method based on TDVP and gradient-descent was developed to identify periodic trajectories in the Floquet Ising model.
@@ Acknowledgements.
We thank Jean-Yves Desaules for useful comments. J.R., A.H. and Z.P. acknowledge support by the Leverhulme Trust Research Leadership Award RL-2019-015 and EPSRC Grant EP/Z533634/1. Statement of compliance with EPSRC policy framework on research data: This publication is theoretical work that does not require supporting research data. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). L.Y. acknowledges support from the Zhejiang Provincial Natural Science Foundation of China (No. LD25A050002), the National Natural Science Foundation of China (Nos. 12247101 and 12375021), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06), and the National Key Research and Development Program of China (No. 2022YFA1404203).
@@ Appendix A Construction of parent Hamiltonians
For a given set of target states $\{|\psi_{n}\rangle\}$, general frameworks exist for constructing parent Hamiltonians, namely, the quantum inverse method [91, 92] and the projective embedding method [63]. For the Type-I scar tower in the spin-1 XY model, we adopt the quantum inverse method, which provides an explicit and interpretable Hamiltonian expression. In contrast, for the Type-II scar tower, the resulting Hamiltonians from the inverse method become overly complicated and obscure the underlying physics. In this case, we instead employ the projective embedding method, which yields a structurally simple and formally well-defined Hamiltonian tailored to the desired subspace.
@@ A.1 Type-I scar tower
The key object in the quantum inverse method is the quantum covariance matrix $C_{T}$ [91, 92], defined on the target subspace:
$(C_{T})_{ab}=\frac{1}{2}\langle\{\hat{h}_{a},\hat{h}_{b}\}\rangle_{T}-\langle\hat{h}_{a}\rangle_{T}\langle\hat{h}_{b}\rangle_{T},$ (20)
where the set of $\{\hat{h}_{a}\}$ is the given list of local operators as the building block for the parent Hamiltonian, and $\langle\cdot\rangle_{T}$ indicates expectation values averaged over all states in the target subspace: $\langle\hat{O}\rangle_{T}=\frac{1}{N}\sum_{n=1}^{N}\langle\psi_{n}|\hat{O}|\psi_{n}\rangle$. For the Type-I scar tower, we start by selecting an operator basis consisting of nearest-neighbor couplings for a spin-1 system. Specifically, we consider two-site interactions generated by the following local operators: $\{\hat{I},\,\hat{S}^{x},\,\hat{S}^{y},\,\hat{S}^{z},\,\hat{P}^{0}\}$, where the operator $\hat{P}^{0}=|0\rangle\langle 0|$ projects a local spin onto the zero magnetization state.
We numerically construct the covariance matrix $C_{T}$ [Eq. (20)] using exact Type-I scar eigenstates as input for a finite system of length $L=8$. Explicitly, the scar eigenstates used in this construction are given by Eq. (9). By computing the numerical null space of $C_{T}$, we identify a set of 9 linearly independent Hamiltonian terms under which the given scar states form an exact degenerate eigenspace. These Hamiltonians take the following explicit forms:
@@
In Ref. [17], the additional perturbation included was simply $\hat{H}_{1}$. In the main text, we focus on a representative Hamiltonian of type $\hat{H}^{l}$ with the specific choice $\hat{V}_{j}=\hat{S}^{x}_{j}$. For a more robust and generic construction yielding a cleaner scar model, we further propose the following random Hermitian operator:
@@
where $\hat{\lambda}^{(n)}$ is the $n$-th Gell-Mann matrix, providing a maximally generic perturbation that eliminates residual structure and promotes optimal convergence of the ScarFinder algorithm.

### v3

@@ V Identifying stable islands in Poincaré sections
In few-body dynamical systems, periodic dynamics and atypical eigenstates can emerge from the coexistence of regular and chaotic regions, resulting from weakly-broken integrability described by the Kolmogorov-Arnold-Moser (KAM) theorem [ArnoldBook, GutzwillerBook]. This phenomenon is known as mixed phase space and it has many physical realizations, including the Chirikov map, the Fermi-Pasta-Ulam system of coupled oscillators, the Lorentz attractor, and many other [stockmann2000quantum]. For a general dynamical system, mixed phase space is diagnosed by recording successive intersections of trajectories with a chosen lower-dimensional transverse surface in phase space – a map known as the Poincaré section [strogatz2001nonlinear]. Periodic trajectories correspond to stationary points of the Poincaré map, while during the chaotic dynamics, the system returns to the same surface at a location that is generally far away from a previous encounter.
Ref. [Michailidis2020] demonstrated that analogues of Poincaré sections can be identified in many-body quantum systems, once their dynamics is projected to a low-dimensional MPS manifold. The resulting TDVP equations of motion within the manifold then serve as an effective mapping of the many-body quantum system onto a few-body classical (non-linear) dynamical system. For the latter, standard tools [strogatz2001nonlinear] can be employed to obtain Poincaré sections. However, such analyses become increasingly cumbersome in higher-dimensional manifolds. A pertinent practical question is: can one directly identify the centers of KAM tori associated with scar dynamics?
From the above description of the Poincaré section, it should be clear that it has many things in common with the ScarFinder algorithm. As discussed in Sec. II, the ScarFinder algorithm can be viewed as generating discrete-time dynamics on the variational manifold. Although the theoretical justification of ScarFinder naturally favors larger projection time steps $\Delta t$, the method remains well-defined in the continuous-time limit $\Delta t\rightarrow 0$. In this limit, ScarFinder dynamics become equivalent to the TDVP evolution equations [Michailidis2020]. This connection implies that, for a given MPS ansatz describing the manifold $\mathcal{M}$, the continuous-time limit $\Delta t\rightarrow 0$ of the ScarFinder algorithm should recover precisely the TDVP dynamics within $\mathcal{M}$. Since any numerical implementation of ScarFinder necessarily uses finite time steps $\Delta t$, such discretizations will inevitably introduce small, systematic deviations from the ideal TDVP trajectory. Surprisingly, rather than causing the dynamics to become chaotic, these deviations accumulate constructively, driving the trajectory towards stable fixed points—precisely the centers of stable islands in the mixed phase space portrait. This motivates the following generalization of the ScarFinder algorithm for exploring Poincaŕe sections.
@@ V.1 ScarFinder algorithm for stable islands in Poincaré sections
Input: An appropriate manifold $\mathcal{M}$, Hamiltonian $\hat{H}$, projection time step $\Delta t$, number of iteration steps $N_{\text{step}}$, Poincaré sampling number $N_{\text{sample}}$.
@@
4:   Evolve $|\psi\rangle$ to $|\psi(\Delta)\rangle\coloneq\exp(-i\hat{H}\Delta t)|\psi\rangle$;
5:    Project $|\psi(\Delta t)\rangle$ back onto $\mathcal{M}$ by maximizing the logarithmic overlap, $\max_{A({\bm{\theta}})}\mathcal{N}[\psi^{\prime},\psi(A(\bm{\theta}))]$, $|\psi^{\prime}\rangle\coloneq\hat{P}_{\mathcal{M}}|\psi(\Delta t)\rangle$;
@@
9:   Update $|\psi\rangle\coloneq|\psi^{\prime}\rangle$.
@@
We note that, in general, the projection step (line 5) should include an energy correction to ensure that the target energy density is preserved. This can be achieved by slightly modifying the projection procedure to incorporate the energy constraint during the maximization. However, for certain models such as the PXP model studied in Sec. V.2 below, this correction turns out to be unnecessary, as the chosen variational manifold inherently preserves the energy.
@@ V.2 Application to PXP mixed phase space
Figure 6: (a),(b) Comparison of Poincaré sections obtained via TDVP and ScarFinder. Trajectories obtained from TDVP (blue points) and ScarFinder (orange points with numbers indicating iteration steps), starting from the same initial condition. The trajectory under ScarFinder spirals rapidly towards the center of the stable island, corresponding exactly to the robust scar trajectory identified as the “*-state” in Ref. [Michailidis2020]. The figure is divided into two panels, (a) and (b), for clarity, showing two distinct regions where the trajectories cross the plane. (c) Our reproduction of the TDVP Poincaré section from Ref. [Michailidis2020], displaying a mixture of stable islands and chaotic regions. (d) ScarFinder directly identifies the centers of stable islands in (c).
Let us illustrate the Poincaré-ScarFinder algorithm on the example considered in Ref. [Michailidis2020]. We assume the PXP model with a simple two-dimensional manifold defined by the following MPS ansatz, first introduced in Ref. [Ho2019]:
@@
Figures 6(a) and 6(b) illustrate the contrast between TDVP and ScarFinder dynamics. While TDVP trajectories trace invariant tori, the finite-step ScarFinder trajectory spirals to a unique attractor, located precisely at the center of a stable island. In Fig. 6(c)-(d), we present additional examples, confirming that ScarFinder systematically identifies the main stable-island centers, provided that a sufficient number of initial conditions are sampled.
@@
For smaller islands, we observe minor discrepancies between the locations of fixed points found by ScarFinder and the island centers identified from TDVP-generated Poincaré sections. To analyze these differences, we extensively sample 640 random initial states within the three-site unit-cell MPS manifold for the PXP model, Eq. (19), and identify a total of 16 distinct fixed-point solutions (within $\theta_{1}\in[0.5\pi,1.5\pi],\ \theta_{3}\in[0,\pi]$) using the ScarFinder algorithm, as shown in Fig.7(a). We find that the fixed points labeled “1” and “2” correspond exactly (within numerical precision) to the “*-state” previously discovered in Ref. [Michailidis2020]. These points coincide precisely with the centers of the largest stable islands identified from the TDVP-generated Poincaré sections.
@@
Figure 7: (a) Distribution of 16 distinct fixed-point solutions identified by ScarFinder from a total of 640 random initial states within the MPS manifold given by Eq. (19). We label the fixed-point sets according to the frequency of getting that solution. The red ellipse highlights two points obtained via different methods: the red star is the optimized state from TDVP-Poincaré section, while point 5 was found by ScarFinder. (b) Comparison of bipartite entanglement entropy dynamics starting from optimized initial states obtained via TDVP Poincaré section [blue line, for the initial state marked by the red star in (a)] and ScarFinder [orange line, for the initial state ”5” in (a)]. (c) Fidelity revival dynamics for the two optimized scar initial states. The state found by ScarFinder exhibits slightly enhanced revivals.
In contrast, the fixed-point set labeled “5” [inside the red ellipse in Fig. 7(a)] identified by ScarFinder is located at $\bm{\theta}_{\mathrm{SF}}{=}(0.928,\,0,\,0.588)\pi$, while the corresponding stable island center from the TDVP-generated Poincaré section is at: $\bm{\theta}_{\mathrm{TDVP}}{=}(0.942,\,0,\,0.5)\pi$. To evaluate the impact of these differences, in Fig. 7(b)-(c) we explicitly compare the dynamics starting from these two initial states. While the dynamical behaviors are qualitatively similar, we observe that the initial state identified by ScarFinder consistently exhibits lower entanglement entropy and noticeably enhanced fidelity revivals.
@@
We thus conclude that the ScarFinder algorithm not only efficiently locates optimal scar states, but in certain cases, even provides better initial conditions compared to states obtained from direct exploration of TDVP-generated mixed-phase spaces. This can be understood by noting that TDVP restricts quantum dynamics to a particular variational ansatz, unlike ScarFinder which utilizes information from exact quantum dynamics during each iteration. The deviation between TDVP and the exact quantum dynamics is characterized using “quantum leakage” [Choi2019]. Since scarred trajectories are typically found in regions of small leakage, there is a strong correspondence between their TDVP trajectory and exact quantum dynamics. However, even in these cases the leakage is non-zero and some relevant information has been lost, therefore it is unsurprising that ScarFinder, can improve upon these results as it does not suffer from any leakage effects.
@@ VI Conclusions
Beyond reproducing the known QMBS phenomenology, our application of ScarFinder to the PXP model gave several intriguing new results. In particular, the observations in Fig. 4 and Fig. 5 hint at an equal energy spacing between (some) scar eigenstates in the thermodynamic limit. Previously, the scar dynamics observed from the $|\mathbb{Z}_{2}\rangle$ state were attributed to an approximately equal energy spacing among these eigenstates [turner2018weak, Choi2019]. However, since the $|\mathbb{Z}_{2}\rangle$ state does not exhibit perfect revivals and exact diagonalization is limited by finite-size effects, it has remained unclear whether the equal energy spacing persists in the thermodynamic limit. Our improved superposition state in the form of an iMPS and the trend in Figs. 4-5 suggest that equal energy spacing is an exact property of some eigenstates in the middle of the spectrum of the PXP model in the thermodynamic limit. This is in line with recent work [ivanov2025exactarealawscareigenstates] which found several exact MPS eigenstates in the spectrum of the PXP model.
Although our discussion primarily focused on 1D systems, the core methodology of ScarFinder generalizes naturally to higher dimensions. In Appendix D, we explore its application to quasi-1D settings, including the PXP model on square and triangular lattices wrapped around a finite cylinder. These results demonstrate the versatility of ScarFinder beyond 1D, and for targeting simultaneously the scar initial states as well as scar eigenstates. For isotropic 2D systems, one could consider extending the approach using projected-entangled-pair state representations [CiracRMP], which would involve a significant computational overhead and we therefore leave it as a future direction.
@@
We emphasize that the utility of ScarFinder is not limited to QMBS models and it can also aid the identification and understanding of other types of atypical dynamics in chaotic models. For example, in the paradigmatic quantum Ising model in the presence of both transverse and longitudinal fields, Ref. [Banuls2011] found certain product states that undergo “weak” thermalization in contrast to the majority of other initial states that strongly thermalize. While these states were subsequently related to low-lying quasiparticle excitations [LinMotrunich2017], their existence is nonetheless surprising given that the model is believed to be “fully” chaotic and obey the ETH. In Appendix E, we show that ScarFinder successfully captures these states and allows to “refine” them by including entanglement. More importantly, ScarFinder allows to search for weakly-thermalizing states, like the ones in Ref. [Banuls2011], systematically and in general models, without prior knowledge about their underlying physics.
We conclude by highlighting several advantages of ScarFinder. In 1D systems, the algorithm is numerically efficient compared to exact diagonalization due to its straightforward implementation using MPS. The latter allows ScarFinder to operate directly in the thermodynamic limit, minimizing finite-size effects and providing direct access to stable, infinite-system scar trajectories. On the other hand, DMRG-based searches have recently been adapted to target low-entanglement excited eigenstates in finite systems [Zhang2023DMRGS]. However, this approach does not automatically yield the initial conditions for probing such eigenstates in experiment. As mentioned above, ScarFinder can identify both scar eigenstates and scar initial states, thereby circumventing the challenge of DMRG-type methods when scar eigenstates possess entanglement beyond area law. We note, however, that some volume-law states, such as the rainbow states from Refs. [PhysRevB.105.L060301, Mohapatra2023], can still be amenable to ScarFinder with a suitably modified unit cell of the lattice, which minimizes entanglement. Additionally, we have demonstrated that ScarFinder directly identifies the centers of stable islands in Poincaré sections, eliminating the need for complicated visualizations of high-dimensional phase spaces. These advantages make ScarFinder a versatile tool for identifying non-thermal behaviors in many-body systems, in particular in the field of quantum simulation and designing long-lived quantum states.
@@
Note Added: During the completion of this work, we became aware of a related study by Petrova et al. [petrova2025findingperiodicorbitsprojected], where a different method based on TDVP and gradient-descent was developed to identify periodic trajectories in the Floquet Ising model.
@@ Acknowledgements.
We thank Jean-Yves Desaules for useful comments. J.R., A.H. and Z.P. acknowledge support by the Leverhulme Trust Research Leadership Award RL-2019-015 and EPSRC Grant EP/Z533634/1. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). L.Y. acknowledges support from the Zhejiang Provincial Natural Science Foundation of China (No. LD25A050002), the National Natural Science Foundation of China (Nos. 12247101 and 12375021), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06), and the National Key Research and Development Program of China (No. 2022YFA1404203).
@@ Appendix A Construction of parent Hamiltonians
For a given set of target states $\{|\psi_{n}\rangle\}$, general frameworks exist for constructing parent Hamiltonians, namely, the quantum inverse method [inverse-method, inverse-method-2] and the projective embedding method [ShiraishiMori]. For the Type-I scar tower in the spin-1 XY model, we adopt the quantum inverse method, which provides an explicit and interpretable Hamiltonian expression. In contrast, for the Type-II scar tower, the resulting Hamiltonians from the inverse method become overly complicated and obscure the underlying physics. In this case, we instead employ the projective embedding method, which yields a structurally simple and formally well-defined Hamiltonian tailored to the desired subspace.
@@ A.1 Type-I scar tower
The key object in the quantum inverse method is the quantum covariance matrix $C_{T}$ [inverse-method, inverse-method-2], defined on the target subspace:
$(C_{\text{target}})_{ab}=\frac{1}{2}\langle[\hat{h}_{a},\hat{h}_{b}]_{+}\rangle_{\text{target}}-\langle\hat{h}_{a}\rangle_{\text{target}}\langle\hat{h}_{b}\rangle_{\text{target}},$ (20)
where $[,]_{+}$ denotes the anticommutator and the set $\{\hat{h}_{a}\}$ is the given list of local operators as the building blocks for the parent Hamiltonian, and $\langle\cdot\rangle_{\text{target}}$ indicates expectation values averaged over all states in the target subspace: $\langle\hat{O}\rangle_{\text{target}}=\frac{1}{N}\sum_{n=1}^{N}\langle\psi_{n}|\hat{O}|\psi_{n}\rangle$. For the Type-I scar tower, we start by selecting an operator basis consisting of nearest-neighbor couplings for a spin-1 system. Specifically, we consider two-site interactions generated by the following local operators: $\{\hat{I},\,\hat{S}^{x},\,\hat{S}^{y},\,\hat{S}^{z},\,\hat{P}^{0}\}$, where the operator $\hat{P}^{0}=|0\rangle\langle 0|$ projects a local spin onto the zero state.
We numerically construct the covariance matrix $C_{T}$ [Eq. (20)] using exact Type-I scar eigenstates as input for a finite system of length $L=8$. Explicitly, the scar eigenstates used in this construction are given by Eq. (9). By computing numerically the null space of $C_{T}$, we identify a set of 9 linearly independent Hamiltonian terms under which the given scar states form an exact degenerate eigenspace. These Hamiltonians take the forms:
@@
In Ref. [Schecter2019], the additional perturbation included was simply $\hat{H}_{1}$ in Eq. (21). In the main text, we focus on a representative Hamiltonian of type $\hat{H}^{l}$ with the specific choice $\hat{V}_{j}=\hat{S}^{x}_{j}$. For a more robust and generic construction, we further propose the following random Hermitian operator:
@@
where $\hat{\lambda}^{(n)}$ is the $n$-th Gell-Mann matrix. This fully eliminates any residual structure and promotes optimal convergence of the ScarFinder algorithm.